How many seven-digit numbers with out repetition of digit are there such that the digit are all non-zero and 5 and 6 do not appear consecutively in either order ?

1
Expert's answer
2022-02-07T12:32:53-0500
  1. Seven-digit numbers that contain no 5 or 6. We have seven digits available (1, 2, 3, 4, 7, 8, 9) to fill seven slots, so there are "7!=5040"
  2. Seven-digit numbers that contain a 5 or 6 but not both. There are two choices for the required digit, seven positions for the required digit, and seven digits available to fill the remaining six slots, so there are "2\u00d77\u00d7^{7}P_6=70560"
  3. Seven-digit numbers that contain both 5 and 6, but not adjacent. To count these, we first count the possible ways we can place the 5 and 6. If we place the 5 at either end, we have five positions available for the 6, which gives us 2×5=10 possibilities. If we place the 5 in one of the five medial position, we have four positions available for the 6 so there are 5×4=20 possibilities. For each of these possibilities, we have seven digits available to fill the remaining five slots. So there are a total of "(10+20)\u00d7^7P_5=75600"

These cases cover the ways of meeting the criteria. Adding the counts together, we get:

"5040+70560+75600=151200"